### v2

# Joint Multimodal Transformer for Emotion Recognition in the Wild

Abstract
Systems for multimodal emotion recognition (MMER) can typically outperform unimodal systems by leveraging the inter- and intra-modal relationships between, e.g., visual, textual, physiological, and auditory modalities. In this paper, an MMER method is proposed that relies on a joint multimodal transformer for fusion with key-based cross-attention. This framework aims to exploit the diverse and complementary nature of different modalities to improve predictive accuracy. Separate backbones capture intra-modal spatiotemporal dependencies within each modality over video sequences. Subsequently, a joint multimodal transformer fusion architecture integrates the individual modality embeddings, allowing the model to capture inter-modal and intra-modal relationships effectively. Extensive experiments [^1] Code: https://github.com/PoloWlg/Joint-Multimodal-Transformer-6th-ABAW. on two challenging expression recognition tasks: (1) dimensional emotion recognition on the Affwild2 dataset (with face and voice), and (2) pain estimation on the Biovid dataset (with face and biosensors), indicate that the proposed method can work effectively with different modalities. Empirical results show that MMER systems with our proposed fusion method allow us to outperform relevant baseline and state-of-the-art methods.
Keywords: Affective Computing, Multi-Modal Fusion, Audio-Visual Fusion, Transformers, Cross-Attention, Valence-Arousal Estimation.
Figure 1: (Top) Illustration of the vanilla multimodal transformer fusion architecture. (Bottom) Multimodal transformer fusion with joint representation (in red) for two input sources, A and B.
1 Introduction
Human-computer interaction is applied in a wide range of real-world scenarios, e.g., health care, IoT, and autonomous driving. Researchers have classified human emotions in different ways, most notably according to discrete categories, ordinal intensity levels, and the valence/arousal circumplex [1]. With the recent advancements in deep learning and sensor technologies, research in affective computing has evolved from lab-controlled to real-world (in the wild) scenarios. In the latter, human emotions are usually expressed over a broader spectrum beyond the six basic categorical expressions - anger, disgust, fear, happy, sad, and surprise [9]. Therefore, there is much interest in analyzing and modeling complex and subtle expressions of emotions in real-world scenarios. For instance, the spectrum of emotions can be formulated as dimensional emotional recognition (ER), where complex human emotions are represented along the axes of arousal (intensity) and valence (positiveness).
Multimodal fusion has been widely explored for the problem of video-based ER in the literature [50, 28]. For instance, audio and visual modalities may provide complementary and redundant information over a video sequence. These relationships must be captured to model the intricacies of human emotions effectively. Furthermore, effectively capturing both the intra-modal temporal dependencies within the audio and visual modalities and the inter-modal association across the audio and visual modalities is crucial in developing an effective AER system[47].
Several methods have been proposed for video-based ER and recurrent networks have been employed to capture the intra-modal temporal dependencies from video sequences [50, 28, 5]. Recently, attention-based methods have been introduced to extract features that are the most relevant to downstream tasks. Cross-attention-based methods have also been [48] employed to capture the inter-modal association between the audio, visual, and other modalities. Lu et al. [35] proposed ViLBERT, which was the seminal work in multimodal co-attention. Since then, many transformer-based cross-attention methods have been proposed [43, 57]. These methods, however, cannot effectively capture the intra-modal temporal dynamics. Further, they specialize in capturing the complementary information among the modalities but do not have the mechanism to capture the redundant information explicitly.
The proposed method introduces the third branch with the joint representation of the multiple modalities, as shown in Figure 1. By incorporating a joint representation branch, the model can access additional contextual information that might not be fully captured by cross-attention alone. This can help improve the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences. This helps in mitigating the sensitivity of cross-attention to noisy inputs and improves the overall performance of the system.
Our main contributions are summarized as follows.
(1) This paper proposes a joint multimodal transformer fusion framework that leverages joint modality representations. It captures inter- and intra-modal information in videos using key-based cross-attention. (2) A joint multimodal transformer multimodal fusion that exploits the redundant and complementary associations among modalities. (3) An extensive set of experiments on the challenging emotion recognition (pain estimation on BioVid and dimension valence-arousal assessment on Affwild2) datasets indicate that our proposed multimodal fusion architecture can outperform relevant baseline and state-of-the-art methods.
2 Related Work in Emotion Recognition
2.1 Multimodal Methods
MMER refers to the integration of multiple sources of information (modalities) to improve the accuracy and robustness of automated emotion recognition systems, at the expense of complexity. These modalities typically include visual, audio, textual, and physiological. The seminal work in multimodal deep learning was proposed by Ngiam et al. [39], where the features from the audio and visual modalities were extracted separately, and then autoencoders and Restricted Boltzmann Machines were used to feature fusion. Tzirakis et al. [55] proposed one of the early approaches for A-V fusion for dimensional emotion recognition, in which the visual features were extracted using a ResNet50 and the audio features were obtained using a 1D convolutional neural network (CNN). The modality-specific features were concatenated and fed to a recurrent net for simultaneous temporal modeling and modality fusion. An empirical study was presented by Juan et al. [42], where the authors studied the impact of fine-tuning multiple layers in a pretrained CNN for the visual modality. A two-stream autoencoder with a long short-term memory (LSTM) network was proposed by Nguyen et al. [40] to jointly learn and compact representative features from the visual and audio modalities. A knowledge distillation-based approach was investigated by Schonevald et al. [50] for visual modality. For the audio modality, spectrograms were obtained and fed to a CNN model, and the two modalities were fused using a recurrent net. A novel self-distillation scheme was put forward by Deng et al. [6] to overcome the problem of noisy labels in a multitasking setting. A two-stream aural visual (TSAV) network was proposed by Kuhnke et al. [28], in which the audio features were extracted using a ResNet18, and the visual features were extracted using a 3D-CNN. The obtained embeddings were fed to a specially designed TSAV network for information fusion.
Pain classification is one of the primary problems in affective computing. Researchers have proposed many multimodal datasets for the pain estimation task. The facial activity descriptors method for pain estimation was introduced by Werner et al. [59]. Dragomir et al. [7] propose a subject-independent method from facial images with a residual learning technique. A Sparse LSTM-based method was proposed by Zhi et al. [65] to solve the problem of vanishing gradients in temporal learning. Morabit et al. [38] proposed a data-efficient image transformer. To process multiscale electrodermal activity signals, a SE-Net-based network was proposed by Lu et al. [36]. Multimodal solutions to fuse the physiological and visual modalities were proposed by Werner et al. [59], Kachele et al. [13], and Zhu et al. [65]. Physiological signals are more discriminative for pain classification than the visual modality.
2.2 Attention-Based and Transformer Methods
Since its inception, attention models have shown extraordinary performance in many applications. These models have been extensively investigated for capturing the inter and intra-modal associations between the audio and visual modalities for tasks like action localization [32], A-V event localization [8], and multimodal emotion recognition [43]. An attention-based fusion mechanism was proposed by Zhang et al. [64], 3D-CNNs and 2D-CNNs were used to extract multi-features in the visual modality, and for the audio modality, a 2D-CNN was used to learn representation from spectrograms. Specialized scoring functions were used to re-weight the audio and visual features. Recently, cross-modal attention has shown promising results because of its ability to model inter-modal relationships. Srinivas et al. [43] explored a transformer network with encoder layers, where cross-modal attention is used to fuse audio and visual features for continuous arousal/valence prediction in the wild. Tzirakis et al. [54] explored the idea of cross-attention in conjunction with self-attention. The authors proposed a transformer-based fusion architecture. Although the methods mentioned above have used cross-modal attention with transformers, they do not have any explicit mechanism to capture semantic relevance between the A-V features, particularly the intra-modal correlations. Zhang et al. [63] proposed a method for A-V fusion using leader-follower attentive fusion for continuous arousal/valence prediction. Attention weights are combined with the encoded visual and audio features. Cross attention presented in Praveen et al. [48] has shown a substantial increase in performance by using cross-correlation across the individual features. In contrast, our proposed method uses key-based cross-attention in multimodal transformers and explores the idea of feeding the joint A-V feature vector. By feeding the joint A-V feature representation, the proposed method effectively captures the inter- and intra-modal relationships simultaneously by interacting across itself and the other modalities.
Huang et al. [11] investigated the idea of multi-head attention in transformer-based fusion architecture, which was further combined with LSTM to capture the high-level representations. Tran et al. [53] proposed a cross-modal transformer architecture that consisted of a multimodal cross-modal attention block, where the Queries were generated from one modality and the key-values were generated from the other modality. Le et al. [31] put forward an end-to-end transformer-based fusion mechanism for multilabel multimodal emotion classification; the model consisted of three parts: i) three backbone networks for visual, audio, and textual feature extractor, ii) a transformer network for information fusion, and iii) classification network. Zhou et al. [66] proposed a transformer-based fusion scheme along with the temporal convolutional network (TCN); the audio and visual features were extracted using pretrained backbones followed by a TCN, the output of TCN was concatenated and fed to a transformer encoder block. A multilayer perceptron (MLP) was then used for the final prediction.
All the aforementioned transformer-based fusion architectures primarily focus on intermodality correlation. In contrast, the proposed method, in addition to modeling the intermodality relationships to capture the complementarity between modalities, we explicitly feed the joint (combined) features to the multimodal transformer to introduce redundancy. By incorporating this third joint representation branch, the proposed model has access to enhanced contextual information that might not be fully captured by cross-attention alone. Doing this improves the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences. Having this third joint representation allows the model to dynamically focus on this newly introduced information in sequences where both modalities are noisy simultaneously. This helps in mitigating the sensitivity of cross-attention to noisy inputs and improves the overall performance of the system.
Figure 2: An overview of the proposed joint multimodal transformer model for A-V fusion. The audio and visual modalities are cross-attended using transformer blocks. The JMT block also takes in the joint representation (shown with red arrows). The output of the cross-attended features is concatenated, and an FC layer is used for valence/arousal prediction.
3 Proposed Approach
The proposed method is a hierarchical fusion mechanism, where the intra-modality features are combined using transformer-based self-attention, and cross-modality features are fused using transformer-based cross-attention. Further, we feed a third joint representation to the joint transformer module (JTM) to enhance robustness. The $K$ (key matrix), $V$ (value matrix), and $Q$ (query matrix) vectors are shared among the six transformer blocks. In the end, the output of these six blocks is again fed to a transformer self-attention block to dynamically weigh the most relevant representations. The final prediction is made using fully connected (FC) layers.
3.1 Modality Specific Feature Extraction
In the first step, modality-specific features are extracted using backbones. The proposed method allows to combine multiple backbones for each modality to improve system robustness. The extracted feature vectors from each backbone are fed to a transformer self-attention block. The combined feature vector is the representation of the particular modality. For example, to capture information about a person’s emotional state, we can use an R(2+1)D CNN pretrained on the Kinetics-400 dataset [29] to extract visual features. For the audio modality, we could extract features using a ResNet18 [10] CNN with a GRU [4].
3.2 Multi-transformer Attention Fusion
We define $\bm{F_{A}}$ as the deep features extracted from backbone A, and $\bm{F_{B}}$ as the deep features extracted from the backbone B in Figure 2. Given $\bm{F_{B}}$ and $\bm{F_{A}}$, the joint feature representation is obtained by concatenating $\bm{F_{B}}$ and $\bm{F_{A}}$ feature vectors:
${\bm{J}}=[{\bm{F}}_{\mathbf{B}};{\bm{F}}_{\mathbf{A}}]\;,$ (1)
[…]
Each representation is then fed to a specific encoder. Our model is composed of three different encoders, one for each type of feature ${\bm{F}}_{\mathbf{B}}$, ${\bm{F}}_{\mathbf{A}}$, and $\bm{J_{AB}}$. Each encoder consists of a multi-head-self-attention (Eq. 2) followed by a fully connected feed-forward network. Residual connection and layer normalization are performed around both of these layers. The key is used to associate a sequence to a key value, the value matrix holds information that is ultimately used to compute the output of the attention mechanism, and the query matrix represents a set of vectors used to query the key-value pairs. The ${K}$, ${V}$, ${Q}$ matrix are calculated this way: ${K={X}{W_{K}}}$, ${Q={X}{W_{Q}}}$, ${V={X}{W_{V}}}$. ${X}$ corresponds to one of the sources of information and ${W_{K}}$, ${W_{Q}}$, ${W_{V}}$ are the weights of the key, query, value matrices respectively. The output values of the self-attention layers are given by:
${Attention(Q,K,V)=softmax({K}{Q^{T}}/{d_{k}})V}$ (2)
With self-attention layers, each encoder focuses independently on important cues related to its respective source of information. Afterward, each encoder embedding is combined by utilizing six cross-modal attention layers where the query matrix Q is shared with the key K and value V matrix of the other source of information. Sharing this matrix between each source of information helps the model add redundancy and complement the visual and audio modalities, thus improving its performance. At the output of each of the six cross-attention modules, the feature vector of dimension 512 is output. These six feature vectors are then stacked to form a sequence, which is then fed to a transformer self-attention block. This block dynamically selects and weighs these feature vectors. The final attended features are fed to an FC layer for final prediction.
The model aims to maximize the Concordance Correlation Coefficient (CCC) [30], which is common in dimensional emotion recognition. To achieve this, we minimize the following loss:
[…]
where ${\rho_{xy}^{2}}$ is the covariance between the predictions and the ground truth, ${\rho_{x}^{2}}$ and ${\rho_{y}^{2}}$ the variance of the prediction and the ground truth, ${\mu_{x}}$ and ${\mu_{y}}$ the mean of prediction and the ground truth.
Table 1: Comparison with the state-of-the-art pain estimation methods on the Biovid Heat Pain Database. The highest score is indicated in bold.
Method Modality CV Scheme Accuracy (%)
Werner et al. [60] ICPR 2014 EDA, ECG, EMG + Video 5-FOLD 80.6
Werner et al. [58] IEEE TAC 2016 Video LOSO 72.4
Kachele et al. [12] IEEE IJSTSP 2016 EDA, ECG, EMG LOSO 82.73
Lopez et al. [33] ACIIW 2017 EDA, ECG 10-FOLD 82.75
Lopez et al. [34] EMBC 2018 EDA LOSO 74.21
Thiam et al. [51] Sensors 2019 EDA LOSO 84.57
Wang et al. [56] EMBC 2020 EDA, ECG, EMG LOSO 83.3
Pouromran et al. [45] PLoS ONE 2021 EDA LOSO 83.3
Thiam et al. [52] Frontiers 2021 EDA, ECG, EMG LOSO 84.25
Phan et al. [44] IEEE Access 2023 EDA, ECG, EMG LOSO 84.8
1D CNN EDA 5-FOLD 77.2
R3D model Video 5-FOLD 72.9
Feature concatenation EDA + Video 5-FOLD 83.5
Vanilla transformer fusion EDA + Video 5-FOLD 87.8
Joint Multimodal Transformer (ours) EDA + Video 5-FOLD 89.1
4 Experimental Methodology
4.1 Datasets
Affwild2
was put forward by [23]. It is one of the largest and most comprehensive datasets for affective computing focused in-the-wild scenarios. The dataset is also a part of the Affective Behavior Analysis in the Wild (ABAW) challenge [2, 22, 21, 17, 16, 3, 15, 27, 26, 20, 18, 25, 24, 19, 61]. The dataset comprises $564$ videos that were collected from YouTube. These videos are labeled for three main affective computing tasks: i) categorical expression recognition, ii) continuous valence arousal prediction, and iii) action unit detection. The dataset comes with a train, validation, and test split with 341, 71, and 152 videos, respectively. The continuous valence/arousal annotation set is used to validate the proposed method.
Biovid Heat Pain Database:
The BioVid Heat Pain Database [60] comprises 87 subjects where heat pain was induced experimentally at the right arm with four different intensities. The data comprises video and depth map video from a Kinect camera, galvanic skin response (EDA), electromyograph (EMG) on trapezius muscle, and electrocardiogram (ECG). The Biovid dataset has various partitions from Part A through E. These partitions differ in the modalities, annotations, and tasks. We use Part A of the dataset and utilize the video and raw EDA data to validate our proposed method.
4.2 Implementation Details
Affwild2:
In the visual modality, cropped and aligned facial images provided with the dataset are used [27]. Black frames (zero pixels) replace missing frames in the visual modality. These images are then fed to a 3D network that takes the input size of 224$\times$224. A clip length of 8 is used, which makes up a sub-sequence of 64 frames. Each sub-sequence contains eight clips. A dropout with a value of 0.8 was used in the linear layers for network regularization. The network was optimized using the stochastic gradient descent (SGD) iterative method with an initial learning rate (LR) of $10^{-3}$. The batch size used for the visual modality was 8. For further generalization, random cropping and random horizontal flips were added as data augmentation. The maximum number of epochs is 50, with early stopping for model selection. In the audio modality, the audio from each video is separated and resampled for 44 kHz. Following the segmentation in the visual modality, small vocal segments are extracted corresponding to the length of the clip. A ResNet18 is used to extract the features from the spectrograms. A Discrete Fourier transform with a length of 1024, a hop length of 10 msec, and a window length of 20 msec is used to obtain spectrograms of each vocal segment. The spectrograms have a resolution of 64$\times$107 pixels, corresponding to a single clip in the visual modality. Other preprocessing of spectrograms include mean and variance normalization, as well as conversion to log-power spectrum. The first convolutional layer in the pretrained ResNet model is adapted to take in the single-channel spectrograms. The learning rate of $1\times 10^{-2}$ is used and Adam optimizer is used to optimize the network. 64 batch size is set for the audio modality.
The audio and visual backbones are frozen to train the A-V fusion network and only train the whole transformer-based fusion model. Each of these backbones outputs deep feature vectors of dimension 512. These features are concatenated (to obtain the joint feature representation, a feature vector of dimension 1024) and then fed to an FC layer to reduce dimensionality to 512. Each of these features is fed to the fusion model, as we can see in Figure 2. At the output of each of the six cross-attention modules, the feature vector of dimension 512 is outputted. These six feature vectors are then stacked to form a sequence, which is then fed to a transformer self-attention block. This block dynamically selects and weighs these feature vectors. The final attended features are fed to a FC layer for final prediction. We perform a grid search to find the optimal learning rate and batch size for our fusion network. Thus, we use a list of learning rates: [$8\times 10^{-4}$, $6\times 10^{-4}$, $3\times 10^{-4}$] and an SGD optimizer to optimize the fusion network. We use a batch size of 32, and the maximum number of epochs is 5, with early stopping for fusion model selection. We select the best model among the learning rates listed before.
Biovid:
In the visual modality, the faces are cropped and aligned using an MTCNN. We apply the frame retention strategy for missing frames where the face is not visible, and the MTCNN cannot capture any frame. Further, to ensure noise-free input to the visual model, we clip the first 2 seconds of the video and the last 0.5 seconds at the end because the subjects show no signs of pain during this duration. The total number of frames is 75. The extracted faces are fed to an R3D model for a visual feature extractor. The batch size is set to 64. The network is separately optimized using the SGD optimizer and the learning rate of $10^{-3}$.
Table 2: Description of the architecture of the custom 1D-CNN for the physiological backbone.
Layer No. of Size of
type Filters Kernel Stride Output
Input - - - $2816$$\times$ $1$
1st Conv $32$ $5$ $2$ $1406$$\times$ $32$
ReLU - - - $1406$$\times$ $32$
Max Pooling - $2$ $703$$\times$ $32$
2nd Conv $64$ $5$ $1$ $699$$\times$ $64$
ReLU - - - $699$$\times$ $64$
Max Pooling - $2$ $349$$\times$ $64$
1st FC - - - $512$
ReLU - - - $512$
2nd FC - - - $2$
For the physiological modality, the EDA is used. The signal is clipped to correspond to the visual modality and fed to a 1D CNN. The architecture of the custom 1D-CNN is shown in Table 2. The CNN outputs a 512-dimensional feature vector. The model is optimized using the SGD optimizer with a learning rate of $10^{-4}$. The batch size for the physiological backbone is set to 1024.
Figure 3: Illustration of the proposed joint multimodal transformer architecture used for the Biovid pain estimation task. The blue branch shows the visual backbone, and the yellow branch is the physiological backbone. The joint representation is shown with a red block. The three feature vectors are fed into the joint transformer block.
For the modality fusion, the two feature vectors and the joint representation are fed to the joint transformer block, as show in Figure 3. The FC layers are removed from both backbones that were added in the backbone training phase. 512-dimensional feature vectors from visual and physiological backbones are obtained and fed to the joint multimodal transformer module. The backbones are frozen, and the joint transformer block is optimized using the ADAM optimizer with a learning rate of $5\times 10^{-6}$, and the batch size is set to 128.
5 Results and Discussion
5.1 Comparison with the State-of-the-Art
Table 1 presents the performance of our fusion model on the Biovid Heat Pain Database. We performed 5-fold cross-validation to pick up the best average fusion model. It can be seen from the table that the proposed model can achieve state-of-the-art performance with multimodal input while using 5-fold cross-validation. The physiological modality is stronger in the Biovid database. Many studies have validated the models on physiological modality. Our empirical results show that the EDA-only accuracy is 77.2%, whereas the visual-only accuracy is 72.9%. The proposed model can improve over unimodal performance and achieves state-of-the-art performance on the Biovid dataset. We also compare it with standard fusion techniques like feature concatenation. For a fair comparison, we keep all the parameters the same. The proposed model improves 6% over simple feature concatenation and 1.3% over a vanilla multimodal transformer i.e. without a joint representation. Figure 4 shows the visualization of the attention weights generated by the joint transformer model. It can be seen that the model gives more weightage to the physiological modality.
Figure 4: Visualization of attention weights for visual and physiological modalities on the Biovid heat pain database. The facial frames are taken 1400 msec each.
Table 3 shows the valence and arousal CCC values on the Affwild2 official validation set and custom-defined folds to increase generalizability. On the official validation set, the proposed method achieves a 0.666 average with 0.717 valence and 0.614 arousal.
Table 3: CCC for valence and arousal of the fusion model trained on different folds of the Affwild2 validation set. Highest scores are indicated in bold.
Validation Set Valence Arousal Average
Official 0.717 0.614 0.666
fold-1 0.705 0.683 0.694
fold-2 0.741 0.623 0.682
fold-3 0.657 0.637 0.647
fold-4 0.760 0.666 0.713
fold-5 0.684 0.629 0.657
Table 4 compares the proposed method with the state-of-the-art on Affwild2 development set. Our method achieves a 0.722 average with 0.76 valence and 0.68 arousal.
Table 4: CCC performance of the proposed and state-of-the-art methods for A-V fusion on the Affwild2 development set.
Method Valence Arousal Mean
Joint Multimodal Transformer (ours) 0.760 0.684 0.722
Situ-RUCAIM3 [37] 0.655 0.709 0.682
Joint Cross- Attention [46] 0.670 0.590 0.630
FlyingPigs [62] 0.559 0.671 0.615
PRL [41] 0.390 0.540 0.465
AU-NO [14] 0.361 0.551 0.456
HSE-NN [49] 0.408 0.477 0.443
Table 5 compares the proposed method with the state-of-the-art on the Affwild2 test set. The proposed method can achieve 0.42 valence and 0.46 arousal. The average of valence and arousal is 0.443. The proposed method significantly improves over the baseline, which is 0.17 average. It is essential to mention here that the other methods that achieve higher performance are due to extensive pertaining, the use of additional modalities like text, and the use of more robust backbones. For a fair comparison, we use a similar setting to Praveen et al. [46], which includes similar pretraining and identical backbones for audio and visual modalities. They used joint cross-attention to fuse the two modalities and achieve a 0.369 average. On the other hand, the proposed model uses a joint multimodal transformer-based fusion method and can achieve a 0.443 average. The proposed method improves 7% over the joint cross-attention-based method.
Table 5: CCC performance of the proposed method and state-of-the-art methods for A-V fusion on the Affwild2 test set.
Method Valence Arousal Mean
Situ-RUCAIM3 [37] 0.606 0.596 0.601
FlyingPigs [62] 0.520 0.602 0.561
PRL [41] 0.450 0.445 0.448
Joint Multimodal Transformer (ours) 0.420 0.467 0.443
HSE-NN [49] 0.417 0.454 0.436
AU-NO [14] 0.418 0.407 0.413
Joint Cross- Attention [46] 0.374 0.363 0.369
Baseline 0.180 0.170 0.175
5.2 Ablations
To see the effectiveness of the joint representation of the proposed method, we performed an ablation study. Table 6 shows the results of the proposed method with and without the joint representation. On the Biovid dataset, the joint multimodal transformer improves by 1.3% over the vanilla multimodal transformer.
Table 6: Performance of the proposed approach compared to the vanilla multimodal transformer on test set. Valence and arousal for the Affwild2 dataset and accuracy for the Biovid dataset. I3D + ResNet18 backbones are used for the Affwild2 dataset and R3D + 1D CNN are used for the BioVid dataset. Default training/validation split is used in the Affwild2 dataset and 5-fold cross validation is performed on the BioVid dataset.
Database Method Accuracy
Affwild2 Vanilla Multimodal Transformer V: 0.432 A: 0.410 Avg: 0.421
Joint Multimodal Transformer V: 0.425 A: 0.450 Avg: 0.438
Biovid Vanilla Multimodal Transformer 87.8
Joint Multimodal Transformer 89.1
Table 7 shows that the proposed model improves accuracy over the vanilla multimodal transformer by 1.8% with a R2D1 and ResNet18 backbone on the Affwild2 dataset. With an I3D and a ResNet18 backbone, results are improved by 1.7%, as shown in Table 8. Notice that the results on the test set are different from Table 5 because we used a different train/validation split.
Table 7: CCC performance of the backbones alone, and combined using baselines and the proposed A-V fusion method on the Affwild2 test set. Experiments are performed on the default training-validation split. R2D1 (visual) and ResNet18 (audio) backbones are used for all cases.
Fusion Model Valence Arousal Mean
R2D1 model only (visual) 0.194 0.310 0.252
ResNet18 model only (audio) 0.273 0.246 0.260
Concat + FC layers 0.320 0.327 0.323
Vanilla Transformer 0.376 0.334 0.355
Joint Multimodal Transformer (ours) 0.366 0.379 0.373
Table 8: CCC performance of the backbones alone, and combined using baselines and the proposed A-V fusion method on the Affwild2 test set. Experiments are performed on the default training-validation split. I3D (visual) and ResNet18 (audio) backbones are used for all cases.
Fusion Model Valence Arousal Mean
I3D model only (visual) 0.336 0.422 0.379
ResNet18 model only (audio) 0.273 0.246 0.260
Concat + FC layers 0.387 0.453 0.420
Vanilla Transformer 0.432 0.410 0.421
Joint Multimodal Transformer (ours) 0.425 0.450 0.438
Joint Cross Attention [46] 0.374 0.363 0.369
6 Conclusion
Multimodal emotion recognition systems outperform their unimodal counterparts, especially in the wild environment. The missing and noisy modality is a prevalent issue with in-the-wild emotion recognition systems. Many attention-based methods have been proposed in the literature to overcome this problem. These methods aim to weigh the modalities dynamically. This paper introduces a joint multimodal transformer for dimensional emotional recognition in the wild. This transformer-based architecture introduces a joint feature representation in order to add more redundancy and complementary between audio and visual data. The proposed model outperforms the state of the art on the Biovid dataset and improves 6% on the Affwild2 dataset. In the future, we aim to introduce more modalities and sophisticated backbones for effective feature extraction.